Given that the volume of gaseous products produces in the unit of dm^3g^-1 at STP in an explosion caused by nitroglycerine is 0.9 times of the volume of gaseous products produced in the same conditions in the explosion caused by picric acid and the heat evolved in the unit of Jg^-1 in the explosion cause by nitroglycerine is 1.9 times of the heat evolved (Jg^-1) in the explosion caused by picric acid.

i) Calculate the power index (PI) for nitroglycerine at STP.

ii) If the sum of the volume of gaseous products produced at STP in the explosion caused by picric acid and nitroglycerine is 1.577 dm^3g^-1 , calculate the volume of gaseous products produced in the unit of dm^3g^-1 at STP in the explosion caused by picric acid and nitroglycerine.
1
Expert's answer
2020-12-03T13:54:21-0500

i) The value for the explosive power is compared with the explosive power of a standard explosive (picric acid) resulting in the power index, shown in the equation below:

"PI = \frac{Q}{Q_p}·\frac{V}{V_p}·100\% = 1.9·0.9 ·100\%= 171\%" .

ii) the sum of the volume of gaseous products produced at STP in the explosion caused by picric acid and nitroglycerine is 1.577 dm3g-1:

"V + V_p = 1.577" ,

while their ratio is:

"\frac{V}{V_p} = 0.9" .

From these two relations, one can make a system and solve it:

"0.9\u00b7V_p + V_p = 1.577"

"V_p = 0.83"

"V = 1.577 - V_p = 1.577 - 0.83 = 0.747" dm3g-1.

Answer: i) the power index (PI) for nitroglycerine at STP is 171%; ii) the volume of gaseous products produced in the unit of dm3g-1 at STP in the explosion caused by picric acid and nitroglycerine are 0.83 and 0.747, respectively.